OpenStudy (anonymous):

Carla's plant was 10 2/3 inches tall. Yolanda's was 8 3/4 inches tall. How much taller was Carla's plant than Yolanda's? A.1 11/12 in. B.2 1/12 in. C.2 5/12 in. D. 2 11/12

OpenStudy (anonymous):

\[10+\frac{2}{3}-8-\frac{3}{4}=\frac{23}{12}=1+\frac{11}{12} \]

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

First, find a common denominator and turn the fractions into improper fractions 10 2/3 =32/3, 8 3/4= 35/4. The LCD is 12, so 32/3= 128/12, and 35/4= 105/12. Now subtract to find the difference: 128/12-105/12= 23/12. Change that to 1 11/12 The height difference is 1 11/12 inches